# Asymptotic behaviors and dynamics of degenerate and mixed solitons for the coupled Hirota system with strong coherent coupling effects

### v1

@@ 4. Asymptotic behaviors of the mixed solitons
$q_{1}[3]=-2{\rm i}e^{-2{\rm i}\eta_{1,R}}\left(\begin{array}[]{cccccc}e^{\theta}&0&-l_{13}^{*}&-l_{14}^{*}&e^{-{\rm i}\left(\eta_{3}-\eta_{1}^{*}\right)}&0\end{array}\right)\varOmega[3]^{-1}\left(\begin{array}[]{c}l_{13}^{*}e^{-\theta}\\ -l_{14}^{*}e^{-\theta}\\ 1\\ 0\\ l_{33}^{*}e^{-{\rm i}\left(\eta_{3}^{*}-\eta_{1}\right)}\\ -l_{34}^{*}e^{-{\rm i}\left(\eta_{3}^{*}-\eta_{1}\right)}\end{array}\right),$ (64h)
$q_{2}[3]=-2{\rm i}e^{-2{\rm i}\eta_{1,R}}\left(\begin{array}[]{cccccc}e^{\theta}&0&-l_{13}^{*}&-l_{14}^{*}&e^{-{\rm i}\left(\eta_{3}-\eta_{1}^{*}\right)}&0\end{array}\right)\varOmega[3]^{-1}\left(\begin{array}[]{c}l_{14}^{*}e^{-\theta}\\ l_{13}^{*}e^{-\theta}\\ 0\\ 1\\ l_{34}^{*}e^{-{\rm i}\left(\eta_{3}^{*}-\eta_{1}\right)}\\ l_{33}^{*}e^{-{\rm i}\left(\eta_{3}^{*}-\eta_{1}\right)}\end{array}\right),$ (64p)
@@
where $\varOmega\left(\Gamma_{3},\Gamma_{s}\right)=-\varOmega\left(\Gamma_{s},\Gamma_{3}\right)^{\dagger}=\frac{1}{\lambda_{s}-\lambda_{3}^{*}}\left(H_{3}^{\dagger}H_{s}e^{{\rm i}\left(\eta_{3}^{*}-\eta_{s}\right)}+Y_{3}^{\dagger}Y_{s}e^{-{\rm i}\left(\eta_{3}^{*}-\eta_{s}\right)}\right)$ $(s=1,2,3)$.
@@ A. Case 1: $l_{13}^{2}+l_{14}^{2}\neq 0$ and $l_{33}^{2}+l_{34}^{2}\neq 0$
$\phi_{1}^{(1)-}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+a_{1}^{-}\right),$
$\phi_{2}^{(1)-}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}+{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}+{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+b_{1}^{-}\right),$
@@
$\phi_{1}^{(2)-}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+a_{2}^{-}\right),$
$\phi_{2}^{(2)-}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}+{\rm i}l_{14}\right)^{*}\chi}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}+{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+b_{2}^{-}\right),$
@@
$\phi_{1}^{(3)-}=2\lambda_{3,I}e^{-2{\rm i}\eta_{3,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}-\lambda_{3}^{*})\left(l_{33}-{\rm i}l_{34}\right)^{*}}{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}^{*}-\lambda_{3})\left|l_{33}-{\rm i}l_{34}\right|}{\rm sech}\left(2\eta_{3,I}-a_{3}^{-}\right),$
$\phi_{2}^{(3)-}=2\lambda_{3,I}e^{-2{\rm i}\eta_{3,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}-\lambda_{3}^{*})\left(l_{33}+{\rm i}l_{34}\right)^{*}}{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}^{*}-\lambda_{3})\left|l_{33}+{\rm i}l_{34}\right|}{\rm sech}\left(2\eta_{3,I}-b_{3}^{-}\right),$
@@
$\phi_{1}^{(2)+}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}-\lambda_{3}^{*})\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+a_{2}^{+}\right),$
$\phi_{2}^{(2)+}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}-\lambda_{3}^{*})\left(l_{13}+{\rm i}l_{14}\right)^{*}\chi}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}+{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+b_{2}^{+}\right),$
@@
$\phi_{1}^{(1)+}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}-\lambda_{3}^{*})\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+a_{1}^{+}\right),$
$\phi_{2}^{(1)+}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}-\lambda_{3}^{*})\left(l_{13}+{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}+{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+b_{1}^{+}\right),$
@@
$\phi_{1}^{(3)+}=2\lambda_{3,I}e^{-2{\rm i}\eta_{3,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}^{*}-\lambda_{3})\left(l_{33}-{\rm i}l_{34}\right)^{*}}{(\lambda_{1}-\lambda_{3})(\lambda_{1}-\lambda_{3}^{*})\left|l_{33}-{\rm i}l_{34}\right|}{\rm sech}\left(2\eta_{3,I}-a_{3}^{+}\right),$
$\phi_{2}^{(3)+}=2\lambda_{3,I}e^{-2{\rm i}\eta_{3,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}^{*}-\lambda_{3})\left(l_{33}+{\rm i}l_{34}\right)^{*}}{(\lambda_{1}-\lambda_{3})(\lambda_{1}-\lambda_{3}^{*})\left|l_{33}+{\rm i}l_{34}\right|}{\rm sech}\left(2\eta_{3,I}-b_{3}^{+}\right),$
@@ B. Case 2: $l_{13}^{2}+l_{14}^{2}\neq 0$ and $l_{33}^{2}+l_{34}^{2}=0$
$\phi_{1B}^{(2)+}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi}{\left|\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+a_{2B}^{+}\right),\quad\phi_{2B}^{(2)+}=\phi_{2}^{(2)+},$
@@
$\phi_{1B}^{(1)+}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+a_{1B}^{+}\right),\quad\phi_{2B}^{(1)+}=\phi_{2}^{(1)+},$
@@ C. Case 3: $l_{13}^{2}+l_{14}^{2}=0$ and $l_{33}^{2}+l_{34}^{2}\neq 0$
$\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)\to\left(\begin{array}[]{c}q_{1}^{{\rm line}-}\\ q_{2}^{{\rm line}-}\end{array}\right)$ (181)
$={\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right)}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right)\right|}\left(\begin{array}[]{c}{\rm sech}\left(\theta-\frac{\hat{\rho}^{-}}{2}\right)\\ -\frac{l_{13}}{l_{14}}{\rm sech}\left(\theta-\frac{\hat{\rho}^{-}}{2}\right)\end{array}\right),$ (184)
@@
$\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)\to\left(\begin{array}[]{c}q_{1}^{{\rm line}+}\\ q_{2}^{{\rm line}+}\end{array}\right)$ (212)
$={\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}-\lambda_{3}^{*})\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right)}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right)\right|}\left(\begin{array}[]{c}{\rm sech}\left(\theta-\frac{\hat{\rho}^{+}}{2}\right)\\ -\frac{l_{13}}{l_{14}}{\rm sech}\left(\theta-\frac{\hat{\rho}^{+}}{2}\right)\end{array}\right),$ (215)
@@
$\phi_{1C}^{(3)+}=2\lambda_{3,I}e^{-2{\rm i}\eta_{3,R}}\frac{\left(l_{33}-{\rm i}l_{34}\right)^{*}}{\left|l_{33}-{\rm i}l_{34}\right|}{\rm sech}\left(2\eta_{3,I}-a_{3C}^{+}\right),\quad\phi_{2C}^{(3)+}=\phi_{2}^{(3)+},$

### v2

@@ 4. Asymptotic behaviors of the mixed solitons
$\hskip-48.36958ptq_{1}[3]=-2{\rm i}e^{-2{\rm i}\eta_{1,R}}\left(\begin{array}[]{cccccc}e^{\theta}&0&-l_{13}^{*}&-l_{14}^{*}&e^{-{\rm i}\left(\eta_{3}-\eta_{1}^{*}\right)}&0\end{array}\right)\varOmega[3]^{-1}\left(\begin{array}[]{c}l_{13}^{*}e^{-\theta}\\ -l_{14}^{*}e^{-\theta}\\ 1\\ 0\\ l_{33}^{*}e^{-{\rm i}\left(\eta_{3}^{*}-\eta_{1}\right)}\\ -l_{34}^{*}e^{-{\rm i}\left(\eta_{3}^{*}-\eta_{1}\right)}\end{array}\right),$ (64h)
$\hskip-48.36958ptq_{2}[3]=-2{\rm i}e^{-2{\rm i}\eta_{1,R}}\left(\begin{array}[]{cccccc}e^{\theta}&0&-l_{13}^{*}&-l_{14}^{*}&e^{-{\rm i}\left(\eta_{3}-\eta_{1}^{*}\right)}&0\end{array}\right)\varOmega[3]^{-1}\left(\begin{array}[]{c}l_{14}^{*}e^{-\theta}\\ l_{13}^{*}e^{-\theta}\\ 0\\ 1\\ l_{34}^{*}e^{-{\rm i}\left(\eta_{3}^{*}-\eta_{1}\right)}\\ l_{33}^{*}e^{-{\rm i}\left(\eta_{3}^{*}-\eta_{1}\right)}\end{array}\right),$ (64p)
@@
[…]
@@ A. Case 1: $l_{13}^{2}+l_{14}^{2}\neq 0$ and $l_{33}^{2}+l_{34}^{2}\neq 0$
$\hskip-28.45274pt\phi_{1}^{(1)-}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+a_{1}^{-}\right),$
$\hskip-28.45274pt\phi_{2}^{(1)-}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}+{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}+{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+b_{1}^{-}\right),$
@@
$\hskip-28.45274pt\phi_{1}^{(2)-}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+a_{2}^{-}\right),$
$\hskip-28.45274pt\phi_{2}^{(2)-}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}+{\rm i}l_{14}\right)^{*}\chi}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}+{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+b_{2}^{-}\right),$
@@
$\hskip-28.45274pt\phi_{1}^{(3)-}=2\lambda_{3,I}e^{-2{\rm i}\eta_{3,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}-\lambda_{3}^{*})\left(l_{33}-{\rm i}l_{34}\right)^{*}}{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}^{*}-\lambda_{3})\left|l_{33}-{\rm i}l_{34}\right|}{\rm sech}\left(2\eta_{3,I}-a_{3}^{-}\right),$
$\hskip-28.45274pt\phi_{2}^{(3)-}=2\lambda_{3,I}e^{-2{\rm i}\eta_{3,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}-\lambda_{3}^{*})\left(l_{33}+{\rm i}l_{34}\right)^{*}}{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}^{*}-\lambda_{3})\left|l_{33}+{\rm i}l_{34}\right|}{\rm sech}\left(2\eta_{3,I}-b_{3}^{-}\right),$
@@
$\hskip-28.45274pt\phi_{1}^{(2)+}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}-\lambda_{3}^{*})\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+a_{2}^{+}\right),$
$\hskip-28.45274pt\phi_{2}^{(2)+}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}-\lambda_{3}^{*})\left(l_{13}+{\rm i}l_{14}\right)^{*}\chi}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}+{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+b_{2}^{+}\right),$
@@
$\hskip-28.45274pt\phi_{1}^{(1)+}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}-\lambda_{3}^{*})\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+a_{1}^{+}\right),$
$\hskip-28.45274pt\phi_{2}^{(1)+}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}-\lambda_{3}^{*})\left(l_{13}+{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(l_{13}+{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+b_{1}^{+}\right),$
@@
$\hskip-28.45274pt\phi_{1}^{(3)+}=2\lambda_{3,I}e^{-2{\rm i}\eta_{3,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}^{*}-\lambda_{3})\left(l_{33}-{\rm i}l_{34}\right)^{*}}{(\lambda_{1}-\lambda_{3})(\lambda_{1}-\lambda_{3}^{*})\left|l_{33}-{\rm i}l_{34}\right|}{\rm sech}\left(2\eta_{3,I}-a_{3}^{+}\right),$
$\hskip-28.45274pt\phi_{2}^{(3)+}=2\lambda_{3,I}e^{-2{\rm i}\eta_{3,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}^{*}-\lambda_{3})\left(l_{33}+{\rm i}l_{34}\right)^{*}}{(\lambda_{1}-\lambda_{3})(\lambda_{1}-\lambda_{3}^{*})\left|l_{33}+{\rm i}l_{34}\right|}{\rm sech}\left(2\eta_{3,I}-b_{3}^{+}\right),$
@@ B. Case 2: $l_{13}^{2}+l_{14}^{2}\neq 0$ and $l_{33}^{2}+l_{34}^{2}=0$
$\hskip-28.45274pt\phi_{1B}^{(2)+}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi}{\left|\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\zeta+a_{2B}^{+}\right),\quad\phi_{2B}^{(2)+}=\phi_{2}^{(2)+},$
@@
$\hskip-28.45274pt\phi_{1B}^{(1)+}=2{\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{\left(l_{13}-{\rm i}l_{14}\right)^{*}\chi^{*}}{\left|\left(l_{13}-{\rm i}l_{14}\right)\chi\right|}{\rm sech}\left(\xi+a_{1B}^{+}\right),\quad\phi_{2B}^{(1)+}=\phi_{2}^{(1)+},$
@@ C. Case 3: $l_{13}^{2}+l_{14}^{2}=0$ and $l_{33}^{2}+l_{34}^{2}\neq 0$
$\hskip-48.36958pt\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)\to\left(\begin{array}[]{c}q_{1}^{{\rm line}-}\\ q_{2}^{{\rm line}-}\end{array}\right)$ (181)
$\hskip-8.5359pt={\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right)}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right)\right|}\left(\begin{array}[]{c}{\rm sech}\left(\theta-\frac{\hat{\rho}^{-}}{2}\right)\\ -\frac{l_{13}}{l_{14}}{\rm sech}\left(\theta-\frac{\hat{\rho}^{-}}{2}\right)\end{array}\right),$ (184)
@@
$\hskip-48.36958pt\left(\begin{array}[]{c}q_{1}\\ q_{2}\end{array}\right)\to\left(\begin{array}[]{c}q_{1}^{{\rm line}+}\\ q_{2}^{{\rm line}+}\end{array}\right)$ (212)
$\hskip-8.5359pt={\rm i}\lambda_{1,I}e^{-2{\rm i}\eta_{1,R}}\frac{(\lambda_{1}^{*}-\lambda_{3}^{*})(\lambda_{1}-\lambda_{3}^{*})\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right)}{\left|(\lambda_{1}-\lambda_{3})(\lambda_{1}^{*}-\lambda_{3})\left(m_{11}-\frac{l_{13}}{l_{14}}m_{12}\right)\right|}\left(\begin{array}[]{c}{\rm sech}\left(\theta-\frac{\hat{\rho}^{+}}{2}\right)\\ -\frac{l_{13}}{l_{14}}{\rm sech}\left(\theta-\frac{\hat{\rho}^{+}}{2}\right)\end{array}\right),$ (215)
@@
$\hskip-28.45274pt\phi_{1C}^{(3)+}=2\lambda_{3,I}e^{-2{\rm i}\eta_{3,R}}\frac{\left(l_{33}-{\rm i}l_{34}\right)^{*}}{\left|l_{33}-{\rm i}l_{34}\right|}{\rm sech}\left(2\eta_{3,I}-a_{3C}^{+}\right),\quad\phi_{2C}^{(3)+}=\phi_{2}^{(3)+},$